2 wires of same dimenssions but ofvdifferent resistivity p1 and p2 if they are joined end to end find the effective resistivity of their combunation

Grade:12

1 Answers

Arun
6 years ago
Dear student
 

R1= ρ1(l/A)

R2= ρ2(l/A)

R= ρ(2l)/A

and Rs = R1+R2

» ρs = (ρ12)/2

 

Regards

Arun (askIITians forum expert)